International Conference on

Geometry of Moduli Spaces and Integrable Systems

RIMS, Kyoto University

September, 3 (Mon) – 7 (Fri), 2001^{1}

Mon  Tues  Wed  Thurs  Fri
9:30–10:30 no talk  Kitaev  Markman  Kostov  Hurtubise
10:45–11:45 no talk  K. Saito Takasaki M.-H. Saito Ohyama
13:00–14:00 Dubrovin Iwasaki  no talk  Takebe  no talk
14:20–15:20  Sakai  Yamada  no talk  Hausel  no talk

15:40–16:40 Boalch Tuda Okumura

no talk Terajima Mano

no talk

September 3, (Monday)

• 13:00–14:00, Boris Dubrovin (SISSA, Trieste): Almost duality for Frobenius manifolds with some applications

• 14:20–15:20, Hidetaka Sakai (Tokyo): Elliptic-difference Painlevé equation

• 15:40–16:40, Philip Boalch (SISSA, Trieste):

Stokes matrices, Poisson Lie groups and Frobenius manifold September 4, (Tuesday)

• 9:30–10:30, Alexander Kitaev (St. Petersburg)

On Algebraic Solutions and Higher Order Transformations for Special Functions of the Isomonodromy Type

• 10:45–11:45 Kyoji Saito (RIMS): Elliptic flat structure and elliptic Lie group

• 13:00–14:00, Katsunori Iwasaki (Kyushu) :

Geometry of isomonodromic deformations

• 14:20–15:20, Yasuhiko Yamada (Kobe) :

Tropical and ultra-discrete Painlevé equations

• 15:40–16:10, Teruhisa Tuda (Tokyo): Folding transformations of the Painlevé equations

1revised on August, 28

• 16:10–16:40 Shoji Okumura (Osaka): The self-dual metrics from Painlevé equations September 5, (Wednesday)

• 9:30–10:30, E. Markman (Amherst):

On the algebraic-geometry of the Classical Dynamical Yang-Baxter Equation

• 10:45–11:45, Kanehisa Takasaki (Kyoto):

Isomonodromic deformations with Calogero degrees of freedom

• No afternoon talk September 6, (Thursday)

• 9:30–10:30, Vladimir Kostov (Nice):

On the Deligne-Simpson problem

• 10:45–11:45, Masa-Hiko Saito (Kobe): Deformation of logarithmic symplectic manifolds and Painlevé equations

• 13:00–14:00, Takashi Takebe (Ochayanomizu) :

Separation of Variables in the Elliptic Gaudin Model

• 14:20–15:20, Tamas Hausel (Berkeley) :

Mirror symmetry, Langlands duality and Hitchin system

• 15:40–16:10 Hitomi Terajima (Kobe):Nodal Curves and Riccati Solutions for Painlevé equations

• 16:10–16:40 Toshiyuki Mano (Kyoto): Differential equations for Hilbert modular

forms of Q(\sqrt{2}) September 7, (Friday)

• 9:30–10:30, Jacques Hurtubise (CRM, Montreal):

The geometry of the Sklyanin bracket

• 10:45–11:45 Yousuke Ohyama (Osaka) :

Special Solutions of Painlevé-Garnier systems

All questions and comments on this conference should be sent to Masa-Hiko Saito. Program will be updated at

http://www.math.kobe-u.ac.jp/HOME/mhsaito/rims2001.html

Department of Mathematics, Faculty of Science Kobe University, Kobe, 657-8501, Japan mhsaito@math.kobe-u.ac.jp fax:+81-(0)78-803-5610 tel:+81-(0)78-803-5614